# FrameMind: Frame-Interleaved Chain-of-Thought for Video Reasoning via Reinforcement Learning

### v1

@@ 2.1 Reasoning in Multimodal LLMs
RL-based post-training has been shown to enhance the reasoning abilities of large (multi)modal models through outcome-driven rewards and group-relative optimization (OpenAI et al., 2024; DeepSeek-AI et al., 2025). A substantial body of work extends this paradigm to multimodal settings by injecting perceptual evidence into the chain of thought: image-side studies expose intermediate spatial cues via visualization or editing (Li et al., 2025), and video-side investigations pursue reinforced video reasoning with multi-task evaluations and trajectory auditing (Feng et al., 2025). Beyond final-answer supervision, structured signals—such as format validity, tool-call correctness, and spatio-temporal alignment—are incorporated to alleviate supervision sparsity and spurious correlations (Zheng et al., 2025; Feng et al., 2025); stability is typically promoted via KL regularization and preference/trajectory auditing (OpenAI et al., 2024; DeepSeek-AI et al., 2025). Notwithstanding these advances, perceptual granularity is often fixed, and attempts to scale RL to hour-level videos reveal unresolved choices concerning the allocation of visual bandwidth under tight computational budgets (Chen et al., 2025).
@@ 3.1 Framework Overview
Problem Formulation. Given a video $V$ and question $Q$, conventional video MLLMs sample a fixed set of frames $F=\{f_{1},f_{2},...,f_{n}\}$ and generate an answer $A$ through a single forward pass: $A=M(F,Q)$.
@@ 3.2 Frame-Interleaved Chain-of-Thought (FiCOT)
At the heart of our agent is the Frame-Interleaved Chain-of-Thought (FiCOT) process, which reformulates video reasoning from a single-step generation into a multi-turn dialogue between reasoning and perception. This grounds the model’s internal monologue in concrete visual evidence. The complete reasoning trajectory is defined as $\tau=\{T_{1},C_{1},E_{1},T_{2},C_{2},E_{2},...,T_{n},A\}$, capturing each step of thought $T_{k}$, action $C_{k}$, observation $E_{k}$, and final answer $A$.
@@ B.1 Detailed Reward Function Setup
    Hard cap at $3$ turns (i.e., if $k=3$ finishes without <answer>, stop).
@@ E.2 Training details
We train a tool-augmented video understanding model based on Qwen/Qwen2.5-VL-7B-Instruct. Optimization uses AdamW (lr $1\times 10^{-6}$, weight decay $1\times 10^{-2}$), no warmup, and gradient clipping at 1.0, with gradient checkpointing enabled. We employ FSDP full sharding (no CPU offload) and tensor parallel size $4$. Batching uses a global batch of $64$ with per-device micro-batches of $2$ for updates and $1$ for experience; padding-free and dynamic batching are enabled. RL follows GRPO with a KL regularizer (low_var_kl, coefficient $10^{-3}$). The visual budget is constrained by min_pixels =50176=50176 and max_pixels =200704=200704, and overlong prompts are kept. For rollouts we use DRFS multi-view with $n=8$ (from lower-res/more-frames to higher-res/fewer-frames), temperature $0.7$, top-$p=0.9$, up to $50$ images per prompt, and target GPU memory utilization $0.7$. Context limits are max_model_len =32768=32768 and max_num_batched_tokens =65536=65536. For validation we adopt conservative settings: temperature $0.01$, top-$p=0.95$, single view $n=1$, and max_frames =64=64. Key settings are summarized in Table E1.

### v2

# framemind: frame-interleaved video reasoning via reinforcement learning

@@ 2.1 Reasoning in Multimodal LLMs
RL-based post-training has been shown to enhance the reasoning abilities of large (multi)modal models through outcome-driven rewards and group-relative optimization (OpenAI et al., 2024; DeepSeek-AI et al., 2025). A substantial body of work extends this paradigm to multimodal settings by injecting perceptual evidence into the chain of thought: image-side studies expose intermediate spatial cues via visualization or editing (Li et al., 2025), representation-level focusing via contrastive attention has also been reported to enhance VLM reasoning (Ge et al., 2025b; a), and video-side investigations pursue reinforced video reasoning with multi-task evaluations and trajectory auditing (Feng et al., 2025). Beyond final-answer supervision, structured signals—such as format validity, tool-call correctness, and spatio-temporal alignment—are incorporated to alleviate supervision sparsity and spurious correlations (Zheng et al., 2025; Feng et al., 2025); stability is typically promoted via KL regularization and preference/trajectory auditing (OpenAI et al., 2024; DeepSeek-AI et al., 2025). Complementary to RL-based methods, recent studies indicate that appropriately designed in-context signals can strengthen reasoning while mitigating overthinking (Ge et al., 2025c); for a broader perspective on prompt and context design practices, see surveys on context engineering (Mei et al., 2025). Notwithstanding these advances, perceptual granularity is often fixed, and attempts to scale RL to hour-level videos reveal unresolved choices concerning the allocation of visual bandwidth under tight computational budgets (Chen et al., 2025).
@@ 3.1 Framework Overview
Problem Formulation. Given a video $V$ and question $Q$, conventional video MLLMs sample a fixed set of frames $F=\{f_{1},f_{2},...,f_{n}\}$ and generate an answer AA through a single forward pass: $A=M(F,Q)$.
@@ 3.2 Frame-Interleaved Chain-of-Thought (FiCOT)
At the heart of our agent is the Frame-Interleaved Chain-of-Thought (FiCOT) process, which reformulates video reasoning from a single-step generation into a multi-turn dialogue between reasoning and perception. This grounds the model’s internal monologue in concrete visual evidence. The complete reasoning trajectory is defined as $\tau=\{T_{1},C_{1},E_{1},T_{2},C_{2},E_{2},...,T_{n},A\}$, capturing each step of thought $T_{k}$, action $C_{k}$, observation $E_{k}$, and final answer AA.
@@ B.1 Detailed Reward Function Setup
    Hard cap at 33 turns (i.e., if $k=3$ finishes without <answer>, stop).
@@ E.2 Training details
We train a tool-augmented video understanding model based on Qwen/Qwen2.5-VL-7B-Instruct. Optimization uses AdamW (lr $1\times 10^{-6}$, weight decay $1\times 10^{-2}$), no warmup, and gradient clipping at 1.0, with gradient checkpointing enabled. We employ FSDP full sharding (no CPU offload) and tensor parallel size 44. Batching uses a global batch of $64$ with per-device micro-batches of 22 for updates and 11 for experience; padding-free and dynamic batching are enabled. RL follows GRPO with a KL regularizer (low_var_kl, coefficient $10^{-3}$). The visual budget is constrained by min_pixels =50176=50176 and max_pixels =200704=200704, and overlong prompts are kept. For rollouts we use DRFS multi-view with $n=8$ (from lower-res/more-frames to higher-res/fewer-frames), temperature $0.7$, top-$p=0.9$, up to $50$ images per prompt, and target GPU memory utilization $0.7$. Context limits are max_model_len =32768=32768 and max_num_batched_tokens =65536=65536. For validation we adopt conservative settings: temperature $0.01$, top-$p=0.95$, single view $n=1$, and max_frames =64=64. Key settings are summarized in Table E1.